Question # 2 (20 marks; 20 min)

Indirect labour (IDL)                                                    $10,600                       $11,500

Indirect materials (IDM)                                                 4,400                           5,000

Total    (MOH)                                                             $15,000                       $16,500

DL(direct labour)                                                         $12,000                       $15,000

MOH =   a   +    b*DL cost

Required:

Determine a     and      b.   (6 marks)

If DL cost is$13,200, how much would be allowed for IDL and IDM, individually?

Hint: Find the linear relationship between MOH and IDM (14 marks)

That is estimate MOH = c + d*IDM
Answer #1

1)Solution :a=$9000 and b=0.5( calculations as given below)

MOH= $ 15,000

MOH= a+ b( Direct Labour Cost)

Therefore $ 15000= a+ b($12000)------equation 1

Similarly $ 16500=a+ b($ 15000)--------equation 2

Taking equation 1 into consideration we get $15000=a+$12000b

i.e a =$15000-$12000b

Substituting the value of “a” as derived above in equation 2 we get,

$ 16500=$15000-$12000b+$15000b

$1500=3000b

Therefore b=0.5

a=$15000-$12000(0.5)=$ 9000

Therefore a=$9000

                                            

2) Amount allowed for Indirect material=$ 4640 and for Indirect labour=$10960(see calculation below)

MOH=c+d(Indirect material)

$15000=c+d($4400)-----equation 3

C=$15000-d($4400)

Substituting this value of c in equation 4

$16500=c+d($5000)-----equation 4

$ 16500=$15000-$4400(d)+$5000(d)

$1500=$600(d)

Therefore d=2.5

C=$15000-$4400(2.5)

Therefore c=$4000

Suppose given direct labour cost=$13200

MOH=a+b(direct labour cost)

MOH=$9000+0.5(direct labour cost)

MOH=$9000+0.5($13200)=$15600

Also given that =c+d(Indirect material)

$15600=$ 4000+2.5( indirect material)

Indirect material=$ 4640

Therefore Indirect labour=$ 15600-$4640=$10960